Project A has an initial cash outflow of $250,000 and is expected to generate $100 thousand a year for the next three years.  Project B has an initial cash outflow of $300 thousand and is expected to generate $125 thousand a year for the next three years.  Assume the discount rate is 8%.

Calculate the NPV and IRR for each project
Answer #1

NPV of Project A =PV of Cash Flows -Initial Investment =100000/(1+8%)+100000/(1+8%)^2+100000/(1+8%)^3-250000
7709.70
NPV of Project B =PV of Cash Flows -Initial Investment =125000/(1+8%)+125000/(1+8%)^2+125000/(1+8%)^3-300000
=22137.12

IRR of Project A using financial calculator
N=3;PMT =100000;PV=-250000;CPT I/Y =9.70%
IRR =9.70%

IRR of Project B using financial calculator
N=3;PMT =125000;PV=-300000;CPT I/Y =12.04%
IRR =12.04%